Hicks spends only shoes and bread. He earns $1,000 and spends $300 of his income on buying new shoes. Suddenly he is promoted and his income is doubled – he now gets $2,000. (a) If his expenditure on shoes becomes $400, the income elasticity of shoes is _______ and shoes are ________ good to Hicks. And bread is _________ good to Hicks. (b) If his expenditure on shoes becomes $700, the income elasticity of shoes is _______ and shoes are ________ good to Hicks. And bread is _________ good to Hicks. (c) If his expenditure on shoes becomes $200, the income elasticity of shoes is _______ and shoes are ________ good to Hicks. And bread is _________ good to Hicks.
Answer #1

Income elasticity = % change in quantity demanded/% change in income

% change in quantity demanded= new expenditure - old expenditure / old expenditure *100

A) income elasticity = 33.33/100= 0.3333

Shoes are normal goods

Bread is normal goods

B) income elasticity = 133.33/100= 1.33

Shoes are normal goods

Bread is normal good

C) income elasticity= -33.33/100= -0.3333

Shoes are inferior good

Bread is a normal good.

If you have any doubt feel free to ask.